the length of a rectangle is 7m less than three times the width. If the length is decreased by 5 and the width by three,the perimeter will be 90m. Find the dimensions of the original rectangle?

Answer by mananth(16946)

the length of a rectangle is 7m less than three times the width.
width =x
length = 3x-7


If the length is decreased by 5 => (3x-7)-5 = 3x-12
and the width by three, => x-3

the perimeter will be 90m.
2( (3x-12)+(x-3)=90
2(4x-15)=90
4x-15 = 45
4x=60
x=15
find length